I am needing some help with these math problems, I can't seem to grasp it.
x - y =4
2x + y = 5

Answer by palanisamy(496)

The given equations are,
x - y =4 ...(1)
2x + y = 5 ...(2)
(1)+(2)=>
3x = 9
x = 9/3
x=3
(2)=> 2*3+y = 5
y = 5-6
y = -1
So the solution is (3,-1)